# Theory uncertainties of the irreducible background to VBF Higgs production

### v1

@@ VII Conclusions
In this study, we have compared theory predictions for the irreducible background to Higgs production in vector boson fusion from gluon-gluon fusion. We began with the prevailing simulation tools used by in ATLAS and CMS, which are based on NNLO precise calculations for the inclusive Higgs boson production process, matched to a parton shower. These calculations achieve LO precision for the final states of interest in the VBF measurements and are therefore below the state of the art. In addition, differences in excess of 20% have previously been observed between different modes of running these simulations. We have pinpointed some of these differences, which have the potential to affect the determination of the CP properties of the Higgs boson.
@@
Overall, we find consistent behavior and agreement of all NLO matched results at the level of 10%. This uncertainty is independent of, and in addition to, the conventional renormalization and factorization scale uncertainties as well as the PDF uncertainties. It is much smaller than the uncertainties observed in previous comparisons of simulation tools employed by the experimental analyses, which indicates that the theory systematics may be poorly understood and overestimated. We note in particular, that we obtain consistent predictions for the dijet azimuthal angle separation, which is of utmost importance for understanding the CP properties of the Higgs boson.
@@ Acknowledgments
Appendix A Impact of the EFT approximation
@@
Appendix B Choice of jet radius
@@
Appendix C Simulations used in LHC experiments
@@
C.1 ATLAS simulations
In the ATLAS experiment, VBF production is simulated with Powheg Box v2 Nason and Oleari (2010); Alioli et al. (2010); Nason (2004); Frixione et al. (2007). Their implementation is based on the corresponding NLO QCD calculations for genuine vector-boson fusion topologies, using the VBF approximation. Quark–antiquark annihilation and interferences between - and -channel contributions are neglected. The renormalization and factorization scales are set to the boson mass, and the PDF4LHC_15nlo PDF set Butterworth et al. (2016) was used. The matrix elements are matched to the “simple” shower algorithm in Pythia 8 Sjöstrand et al. (2015), which employs the AZNLO tune ATLAS Collaboration (2014). The dipole-recoil strategy was utilized for the parton shower. The decays of bottom and charm hadrons are implemented by EVTGEN Lange (2001). The QCD scales and are varied independently by factors of and . The prediction from the Powheg Box sample was normalized to the NNLO cross-section in QCD using the VBF approximation Ciccolini et al. (2007, 2008); Bolzoni et al. (2010). Relative NLO electroweak corrections are also included for the - and -channel contributions considered in the VBF approximation. Alternative predictions are used to estimate the theoretical uncertainty due to variations in parton showering, estimated by showering the same LHE files with the Herwig 7 angular-ordered shower Bellm et al. (2017), using the H7UE set of tuned parameters.
@@ C.1 ATLAS simulations
Multiple ATLAS VBF measurements, such as those from the Aad et al. (2023), have leading systematic uncertainties related to Higgs modeling. Figure 5b in Aad et al. (2023) shows that parton-shower uncertainties are the leading sources, exceeding even experimental uncertainties for jet calibration. This uncertainty is estimated by taking the envelope of the differences between the Powheg Box + Pythia and Powheg Box + Herwig predictions and can reach up to 20%, increasing with jet multiplicities Aad et al. (2023). The differences are typically on yields affecting both VBF and ggF backgrounds. They appear in analyses as differences in acceptances, efficiencies, or expected yields. A sub-dominant uncertainty comes from QCD variations of the ggF background in the VBF-enriched region, as the QCD accuracy of the NNLO Monte Carlo deteriorates when requiring two jets.
C.2 CMS simulations
@@
The CMS Collaboration simulates the VBF Higgs production process at NLO QCD using the Powheg Box v2 Nason and Oleari (2010); Alioli et al. (2010); Nason (2004); Frixione et al. (2007) generator. The dipole-recoil option of the “simple shower” algorithm in Pythia 8 Sjöstrand et al. (2015) is used to obtain a faithful description of the initial-final color flow that takes into account the color connection between the incoming and the outgoing partons. In certain channels, such as the boosted Higgs to channel HIG (2023), the VBF samples are further reweighted to account for NNLO QCD corrections to the spectrum and N3LO QCD corrections to the inclusive cross section Cacciari et al. (2015); Dreyer and Karlberg (2016).
@@
In contrast, the ggF Higgs boson production process is simulated using the HJ-MiNLO event generator within Powheg Box Hamilton et al. (2015, 2013a), with the Higgs boson mass set to , including finite top-quark mass effects. The ggF production process is further simulated with the Powheg Box v2 event generator at NLO in QCD.
@@
For channels such as the Sirunyan et al. (2021a), the kinematic features of events produced in ggF with two associated jets are considered. The MiNLO program at NLO in QCD is utilized for evaluating systematic uncertainties related to the modeling of two associated jets and contamination in the VBF measurement. The two-jet final state is produced with leading order accuracy.
@@
For the decay Sirunyan et al. (2021b, c), the gluon-gluon fusion (ggF) events are simulated at NLO QCD using Powheg Box v2 and additionally reweighted to match the NNLOPS predictions in the distributions of and . The ggF sample is normalized to N3LO QCD accuracy and NLO electroweak accuracy.
@@
In the Hayrapetyan et al. (2024) and Sirunyan et al. (2019), the VBF signal is generated using Powheg Box v2 at NLO QCD accuracy and matched to the “simple shower” in Pythia 8, using the dipole-recoil option. Alternatively, the Powheg Box matrix-element generator interfaced with Herwig 7 is sometimes used for fragmentation and hadronization. This alternative configuration helps to assess the systematic uncertainty related to the choice of showering and hadronization model.
@@
For parton showering and hadronization, most of the generated samples are interfaced with Pythia 8 using the CP5 tune for the underlying-event description, and the NNPDF3.1 set at NNLO QCD accuracy is used for all processes. For Herwig 7, the parton shower relies on the so called CH (CMS Herwig) tunes Sirunyan et al. (2021d), which are based on the NNPDF3.1 NNLO QCD PDF set.
Appendix D Simulation setups and event samples

### v2

@@ VII Conclusions
In this study, we have compared theory predictions for the irreducible background to Higgs production in vector boson fusion from gluon-gluon fusion. We began with the prevailing simulation tools used by ATLAS and CMS, which are based on NNLO precise calculations for the inclusive Higgs boson production process, matched to a parton shower. These calculations achieve LO precision for the final states of interest in the VBF measurements and are therefore below the state of the art. In addition, differences in excess of 20% have previously been observed between different modes of running these simulations. We have pinpointed some of the differences, which have the potential to affect the determination of the CP properties of the Higgs boson.
@@
Overall, we find consistent behavior and agreement of all NLO matched results at the level of 10%. This uncertainty is independent of, and in addition to, the conventional renormalization and factorization scale uncertainties as well as the PDF uncertainties. It is much smaller than the uncertainties observed in previous comparisons of simulation tools employed by the experimental analyses, which indicates that the theory systematics may be poorly understood and overestimated. We note in particular, that we obtain consistent predictions for the dijet azimuthal angle separation, which is important for understanding the CP properties of the Higgs boson.
Our recommendation is therefore to rely on any of the NLO accurate predictions by default, and to gauge the uncertainty of the result by varying among the other NLO accurate setups presented in this study.
@@ Acknowledgments
Appendix A Additional figures
Here, we collect two additional figures not shown in the main text. Figure 14, shows the normalization of the inclusive and NLO+PS predictions to the NLO Higgs rapidity spectrum.
Appendix B Powheg Box predictions with MiNLO scale setting
As an alternative to the Born-suppression method, the process in Powheg Box can be regulated via the MiNLO prescription. As MiNLO both introduces Sudakov effects and changes the scale setting of the hard process, it does not agree with the fixed-order benchmarks set by the other event generators. As such, it is not sensible to include Powheg Box MiNLO predictions in the comparisons shown in the main text. Here, we collect figures comparing Powheg Box + Herwig and Powheg Box + Pythia, including the default shower and the dipole-recoil option, in the MiNLO prescription at hadron level with MPI. The results are shown in Fig. 15.
Appendix C Impact of the EFT approximation
@@
Appendix D Choice of jet radius
@@
Appendix E Simulations used in LHC experiments
@@
E.1 ATLAS simulations
In the ATLAS experiment, VBF production is simulated with Powheg Box v2 Nason and Oleari (2010); Alioli et al. (2010); Nason (2004); Frixione et al. (2007). Their implementation is based on the corresponding NLO QCD calculations for genuine vector-boson fusion topologies, using the VBF approximation. Quark–antiquark annihilation and interferences between - and -channel contributions are neglected. The renormalization and factorization scales are set to the boson mass, and the PDF4LHC_15nlo PDF set Butterworth et al. (2016) was used. The matrix elements are matched to the “simple” shower algorithm in Pythia 8 Sjöstrand et al. (2015), which employs the AZNLO tune Aad et al. (2014). The dipole-recoil strategy was utilized for the parton shower. The decays of bottom and charm hadrons are implemented by EVTGEN Lange (2001). The QCD scales and are varied independently by factors of and . The prediction from the Powheg Box sample was normalized to the NNLO cross-section in QCD using the VBF approximation Ciccolini et al. (2007, 2008); Bolzoni et al. (2010). Relative NLO electroweak corrections are also included for the - and -channel contributions considered in the VBF approximation. Alternative predictions are used to estimate the theoretical uncertainty due to variations in parton showering, estimated by showering the same LHE files with the Herwig 7 angular-ordered shower Bellm et al. (2017), using the H7UE set of tuned parameters.
@@ E.1 ATLAS simulations
Multiple ATLAS VBF measurements, such as those from the Aad et al. (2023b), have leading systematic uncertainties related to Higgs modeling. Figure 5b in Aad et al. (2023b) shows that parton-shower uncertainties are the leading sources, exceeding even experimental uncertainties for jet calibration. This uncertainty is estimated by taking the envelope of the differences between the Powheg Box + Pythia and Powheg Box + Herwig predictions and can reach up to 20%, increasing with jet multiplicities Aad et al. (2023b). The differences are typically on yields affecting both VBF and ggF backgrounds. They appear in analyses as differences in acceptances, efficiencies, or expected yields. A sub-dominant uncertainty comes from QCD variations of the ggF background in the VBF-enriched region, as the QCD accuracy of the NNLO Monte Carlo deteriorates when requiring two jets.
E.2 CMS simulations
@@
The CMS Collaboration simulates the VBF Higgs production process at NLO QCD using the Powheg Box v2 Nason and Oleari (2010); Alioli et al. (2010); Nason (2004); Frixione et al. (2007) generator. The dipole-recoil option of the “simple shower” algorithm in Pythia 8 Sjöstrand et al. (2015) is used to obtain a faithful description of the initial-final color flow that takes into account the color connection between the incoming and the outgoing partons. In certain channels, such as the boosted Higgs to channel CMS collaboration (2023), the VBF samples are further reweighted to account for NNLO QCD corrections to the spectrum and N3LO QCD corrections to the inclusive cross section Cacciari et al. (2015); Dreyer and Karlberg (2016).
@@
In contrast, the ggF Higgs boson production process is simulated using the HJ-MiNLO event generator within Powheg Box Hamilton et al. (2015, 2013a), with the Higgs boson mass set to , including finite top-quark mass effects. Here, Powheg Box v2 provides the NLO QCD matrix-element calculation matched to a parton shower, while HJ-MiNLO specifies the ggF jet setup and MiNLO treatment.
@@
For channels such as the Sirunyan et al. (2021c), the kinematic features of events produced in ggF with two associated jets are considered. The MiNLO program at NLO in QCD is utilized for evaluating systematic uncertainties related to the modeling of two associated jets and contamination in the VBF measurement. The two-jet final state is produced with leading order accuracy.
@@
For the decay Sirunyan et al. (2021d, e), the gluon-gluon fusion (ggF) events are simulated at NLO QCD using Powheg Box v2 and additionally reweighted to match the NNLOPS predictions in the distributions of and . The ggF sample is normalized to N3LO QCD accuracy and NLO electroweak accuracy.
@@
In the Hayrapetyan et al. (2024b) and Sirunyan et al. (2019b), the VBF signal is generated using Powheg Box v2 at NLO QCD accuracy and matched to the “simple shower” in Pythia 8, using the dipole-recoil option. Alternatively, the Powheg Box matrix-element generator interfaced with Herwig 7 is sometimes used for fragmentation and hadronization. This alternative configuration helps to assess the systematic uncertainty related to the choice of showering and hadronization model.
@@
For parton showering and hadronization, most of the generated samples are interfaced with Pythia 8 using the CP5 tune for the underlying-event description, and the NNPDF3.1 set at NNLO QCD accuracy is used for all processes. For Herwig 7, the parton shower relies on the so called CH (CMS Herwig) tunes Sirunyan et al. (2021f), which are based on the NNPDF3.1 NNLO QCD PDF set.
Appendix F Simulation setups and event samples